Members of the Ski Club contributed equally to obtain $1800 for a holiday trip. When 6 members found that they could not go, their contributions were refunded and each remaining member then had to pay $10 more to raise the $1800. How many went on the trip?
Thank you.

Answer by stanbon(75887)

Let number going originally be "x".
Each must have paid 1800/x
If six dropped out that would be 6(1800/x)
The remaining group numbers "x-6".
EQUATION:
1800-6(1800/x)+(x-6)10=1800
6(1800/x)=10(x-6)
10800=10x^2-60x
Divide by 10 to get:
x^2-6x-1080=0
(x-36)(x+30)=0
x=36
There were originally 36 in the group.
Cheers,
Stan H.